ABSOLUTE MEASUREMENTS OF NEAR-INFRARED CO_{2} TRANSITION FREQUENCIES AT THE kHz-LEVEL

\underline{DAVID A. LONG}, GAR-WING TRUONG, JOSEPH T. HODGES, Material Measurement Laboratory, National Institute of Standards and Technology, 100 Bureau Drive, Gaithersburg, MD 20899, USA; and CHARLES E. MILLER, Jet Propulsion Laboratory, California Institute of Technology, 4800 Oak Grove Drive, Pasadena, CA 91109, USA.

Measurements were made of Doppler broadened CO_{2} transitions in the (30013) ⟵ (00001) band near 1.6 μm using frequency-stabilized cavity ring-down spectroscopy (FS-CRDS). These absolute measurements were directly linked to a cesium atomic clock by the use of an octave-spanning, self-referenced optical frequency comb. Combined standard uncertainties as low as 18 kHz were achieved for these weak transitions. A global fit was then performed which included these measurements as well as an ensemble of absolute mid-infrared and far-infrared data. The resulting spectroscopic parameters provide secondary frequency standards over a wide spectral region and should benefit atmospheric remote sensing missions.